OpenStudy (anonymous):

Dosha is creating a new dessert that has two layers shaped like cones. The inner cone is frozen ice cream and has a diameter of 12 cm and a height of 6 cm. The outer layer is a thin wafer shell, like an upside-down ice cream cone, with a height of 15 cm and the same diameter as the inner layer. Dosha will inject a cream filling into the space. What is the volume of the cream filling? Use 3.14 to approximate pi and express your final answer in hundredths.

OpenStudy (anonymous):

\[Volume of cone = \frac{ 1 }{ 3 }\pi*r^2h\]

OpenStudy (anonymous):

Cone 1: Radius :6 cm Height :6 cm Cone 2 Radius :6 cm Height : 15cm

OpenStudy (anonymous):

cone 1= 226.08? cone 2=565.2?

OpenStudy (anonymous):

Cone 1 r = 6 cm h = 6 cm r^2 = r*r = 6*6 = 36 r^2h = r*r*h = 36*6 = 216 Volume of cone = pi*r^2*h/3 = 3.14*216/3

OpenStudy (anonymous):

3.14*72 = 226.08 cm Sorry I was wrong.

OpenStudy (anonymous):

Lemme check the second one

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

r^2 = 6*6 = 36 36*h = 36*!5 = 540 r^h/3 = 540/3 = 180 pi*180 = 3.14*180 = 565.2 cm :)

OpenStudy (anonymous):

do i subtract them now?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

i got 339.12

OpenStudy (anonymous):

Yes, That's should be correct.

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

No problem :)